OpenStudy (anonymous):

HELP :((( Explain how to convert f(x) into the general, vertex form of the equation. Use complete sentences. You may use the f(x) you created in question 1 as an example..... The F(X) example i used previously was f(x)=x^2-4

OpenStudy (anonymous):

not much to do here since it is both in general form as \[f(x)=ax^2+bx+c\] with \(a=1,b=0,c=-4\) and also vertex form as \[f(x)=a(x-h)^2+k\] with \(a=1,h=0, k=-4\)

OpenStudy (anonymous):

Thank you so much, is that all?